solve the following system of equation by the method of your choice.
3x=4y-43
4x-29=-7y

Answer by stanbon(75887)

solve the following system of equation by the method of your choice.
3x=4y-43
4x-29=-7y
-----
Rearrange::
3x - 4y = -43
4x + 7y = 29
-----
Modify for elimination:
12x - 16y = 4*-43
12x + 21y = 3*29
-------------------
Subract and solve for "y"::
37y = 259
y = 7
----
Solve for "x"::
3x-4y = -43
3x-4*7 = -43
3x - 28 = -43
3x = -15
x = -5
------------
Cheers,
Stan H.
------------